# Monocular Lane Detection Based on Deep Learning: A Survey

### v1

@@ VII Future Direction
$\bullet$ General and Unified Lane Modeling. Effectively modeling arbitrarily shaped lanes without compromising efficiency remains a significant challenge. When the scenario extends from a monocular camera’s front view to multi-camera surround views, the presence of numerous U-shaped or nearly horizontal lanes becomes common. In such cases, modeling approaches that heavily rely on prior knowledge, such as grids-based modeling for row-wise classification, or line anchor-based modeling which learn the longitudinal equidistant offset points, are unsuitable. In contrast, mask-based modeling methods are more reliable despite their lower performance and efficiency. The existing vectorized map element modeling methods [170, 148] provide valuable guidance on this issue. Recently, Lane2Seq [141] unified 2D lane detection through sequence generation. This approach also serves as a promising direction for subsequent studies, although its efficiency still requires further improvement.
@@
$\bullet$ Multi-modal Lane Detection. In recent years, LiDAR-based lane detection benchmarks and methods [178, 179, 180, 181, 182, 183, 184] gain attention as another minority approach. Although the 3D information can be directly offered by LiDAR, its shorter perception range and high cost have made camera-based methods more prevalent. However, LiDAR provides the advantages by remaining unaffected by lighting changes and delivering accurate depth information, effectively compensating for the limitations of cameras. The integration of LiDAR and camera data demonstrates significant effectiveness in enhancing performance, which are widely validated in the domain of 3D perception [185, 186, 187, 188]. However, there is currently a lack of dedicated research focusing on multi-modal lane detection [189, 190, 191].
@@
$\bullet$ Label and Model Efficient Lane Detection. The existing lane detection methods mainly focus on supervised learning, which need a lot of annotations for training, which leads to huge manual costs. Thus, developing annotation-efficient lane detection algorithms is necessary. WS-3D-Lane [192] uses 2D lane labels to weakly supervise 3D lane detection, which is valuable for research and mass production. Futhermore, unsupervised lane detection [193, 194, 195, 196] is also a promising direction, although the related works are less.
@@
$\bullet$ Lane Detection in End-to-End Autonomous Driving. The CVPR Best Paper, UniAD [197], generate significant interest [198, 199, 200, 201, 202] in both academia and industry regarding the development of end-to-end autonomous driving systems. Unlike the conventional modular architecture of ”perception, prediction, and planning,” end-to-end autonomous driving directly outputs vehicle motion planning results from sensor data in a fully differentiable manner. Within this framework, lane detection no longer outputs explicit lane coordinate values but instead functions as a module providing intermediate representations of lanes. However, this approach often encounters challenges such as limited interpretability and inadequate generalization, particularly in complex road scenarios, including curved roads and multi-lane switching. Future research can explore hybrid architectures that incorporate specific lane detection outputs—such as lane centerlines, lane width, and curvature—as prior knowledge into intermediate representations within end-to-end models. This integration can enable the network to better capture the structural characteristics of roads and ensure the preservation of critical lane-level information in the decision-making process.
@@
$\bullet$ Visual Reasoning for Lane Detection. The advent of large language models (LLMs) and vision-language Models (VLMs) unlock significant potential for multimodal artificial intelligence systems to perceive the real world, make decisions, and control tools with a capability akin to human cognition. recently, LLMs and VLMs are integrated into autonomous driving systems, primarily focusing on visual reasoning related to dynamic obstacles, the generation of future trajectories, and the detailed control signals of ego-vehicles [203, 204, 205]. In contrast, relatively few studies explore visual inference concerning static obstacles, such as lanes. Fortunately, a new benchmark [206] specifically designed for large-scale visual reasoning in understanding maps and traffic scenarios is emerged. By training on extensive traffic scene data, the models an derive insights from complex multimodal driving resources, including map data, traffic regulations, and incident reports. This enables them to enhance vehicle navigation and planning with safety and efficiency parameters, while also adapting to dynamic road conditions with an understanding that closely resembles human intuition.
@@ VIII Conclusion
This survey comprehensively reviews the latest progress in monocular lane detection based on deep learning, covering both 2D and 3D lane detection methods in recent years. Four core design in lane detection algorithms through theoretical analysis and experimental evaluation are identified: (1) Task paradigm, focusing on lane instance-level discrimination; (2) Lane modeling, representing lanes as a set of learnable parameters in the neural network; (3) Global information supplementation, enhancing the detection of obscured lanes; (4) Perspective effect elimination, providing 3D lanes usable for downstream applications. From these perspectives, this paper presents a comprehensive overview of existing methods, encompassing both the increasingly mature 2D lane detection approaches and the developing 3D lane detection works. In addition, this article also reviews extended work on monocular lane detection to help readers have a more comprehensive understanding of the development process of lane detection. Finally, the future research directions for lane detection are point out.
@@
where $C_{clip},S_{clip}$ are the number of correct points and the number of ground truth points of a image respectively.
@@
where $R$, $T$ refers to a rotation matrix and a translation matrix respectively.
The image coordinate system is employed to represent the perspective projection from the camera coordinate system onto the image plane. When the camera distortion are disregarded, the relationship between a 3D point and its image plane projection can be simplified using a pinhole model. The image coordinates $(x_{i},y_{i})$ are determined by Eqn. 15:
@@
where $f$ represents the focal length of camera.

### v2

@@ VII Future Direction
$\bullet$ General and Unified Lane Modeling. Effectively modeling arbitrarily shaped lanes without compromising efficiency remains a significant challenge. When the scenario extends from a monocular camera’s front view to multi-camera surround views, the presence of numerous U-shaped or nearly horizontal lanes becomes common. In such cases, modeling approaches that heavily rely on prior knowledge, such as grids-based modeling for row-wise classification, or line anchor-based modeling which learns the longitudinal equidistant offset points, are unsuitable. In contrast, mask-based modeling methods are more reliable despite their lower performance and efficiency. The existing vectorized map element modeling methods [170, 148] provide valuable guidance on this issue. Recently, Lane2Seq [141] unified 2D lane detection through sequence generation. This approach also serves as a promising direction for subsequent studies, although its efficiency still requires further improvement.
@@
$\bullet$ Multi-modal Lane Detection. In recent years, LiDAR-based lane detection benchmarks and methods [178, 179, 180, 181, 182, 183, 184] gain attention as another minority approach. Although the 3D information can be directly offered by LiDAR, its shorter perception range and high cost have made camera-based methods more prevalent. However, LiDAR provides the advantages by remaining unaffected by lighting changes and delivering accurate depth information, effectively compensating for the limitations of cameras. The integration of LiDAR and camera data demonstrates significant effectiveness in enhancing performance, which is widely validated in the domain of 3D perception [185, 186, 187, 188]. However, there is currently a lack of dedicated research focusing on multi-modal lane detection [189, 190, 191].
@@
$\bullet$ Label and Model Efficient Lane Detection. The existing lane detection methods mainly focus on supervised learning, which needs a lot of annotations for training, which leads to huge manual costs. Thus, developing annotation-efficient lane detection algorithms is necessary. WS-3D-Lane [192] uses 2D lane labels to weakly supervise 3D lane detection, which is valuable for research and mass production. Furthermore, unsupervised lane detection [193, 194, 195, 196] is also a promising direction, although the related works are less.
@@
$\bullet$ Lane Detection in End-to-End Autonomous Driving. The CVPR Best Paper, UniAD [197], generates significant interest [198, 199, 200, 201, 202] in both academia and industry regarding the development of end-to-end autonomous driving systems. Unlike the conventional modular architecture of ”perception, prediction, and planning,” end-to-end autonomous driving directly outputs vehicle motion planning results from sensor data in a fully differentiable manner. Within this framework, lane detection no longer outputs explicit lane coordinate values but instead functions as a module providing intermediate representations of lanes. However, this approach often encounters challenges such as limited interpretability and inadequate generalization, particularly in complex road scenarios, including curved roads and multi-lane switching. Future research can explore hybrid architectures that incorporate specific lane detection outputs—such as lane centerlines, lane width, and curvature—as prior knowledge into intermediate representations within end-to-end models. This integration can enable the network to better capture the structural characteristics of roads and ensure the preservation of critical lane-level information in the decision-making process.
@@
$\bullet$ Visual Reasoning for Lane Detection. The advent of large language models (LLMs) and vision-language Models (VLMs) unlock significant potential for multimodal artificial intelligence systems to perceive the real world, make decisions, and control tools with a capability akin to human cognition. Recently, LLMs and VLMs are integrated into autonomous driving systems, primarily focusing on visual reasoning related to dynamic obstacles, the generation of future trajectories, and the detailed control signals of ego-vehicles [203, 204, 205]. In contrast, relatively few studies explore visual inference concerning static obstacles, such as lanes. Fortunately, a new benchmark [206] specifically designed for large-scale visual reasoning in understanding maps and traffic scenarios is emerged. By training on extensive traffic scene data, the models can derive insights from complex multimodal driving resources, including map data, traffic regulations, and incident reports. This enables them to enhance vehicle navigation and planning with safety and efficiency parameters, while also adapting to dynamic road conditions with an understanding that closely resembles human intuition.
@@ VIII Conclusion
This survey comprehensively reviews the latest progress in monocular lane detection based on deep learning, covering both 2D and 3D lane detection methods in recent years. Four core designs in lane detection algorithms through theoretical analysis and experimental evaluation are identified: (1) Task paradigm, focusing on lane instance-level discrimination; (2) Lane modeling, representing lanes as a set of learnable parameters in the neural network; (3) Global information supplementation, enhancing the detection of obscure lanes; (4) Perspective effect elimination, providing 3D lanes usable for downstream applications. From these perspectives, this paper presents a comprehensive overview of existing methods, encompassing both the increasingly mature 2D lane detection approaches and the developing 3D lane detection works. In addition, this article also reviews extended work on monocular lane detection to help readers have a more comprehensive understanding of the development process of lane detection. Finally, the future research directions for lane detection are pointed out.
@@
where $C_{clip},S_{clip}$ are the number of correct points and the number of ground truth points of an image respectively.
@@
where $R$, $T$ refer to a rotation matrix and a translation matrix respectively.
The image coordinate system is employed to represent the perspective projection from the camera coordinate system onto the image plane. When the camera distortion is disregarded, the relationship between a 3D point and its image plane projection can be simplified using a pinhole model. The image coordinates $(x_{i},y_{i})$ are determined by Eqn. 15:
@@
where $f$ represents the focal length of the camera.